# Artificial Intelligence LLM Large Language Model

### v2

@@ 2.3.2. Customizations and Hypotheses
Customizations of LLMaaS chatbots that feature the corporate design in the user interface could serve as a salient trust cue for users, increasing the perceived trustworthiness of the system in two dimensions: (a) purpose and (b) process. The perceived purpose of the system may appear more trustworthy when the customizations incorporate university branding (see Figure 2, Label 1) [^1] Figure 2 shows a replica of the LLMaaS chatbot’s user interface, with customizations marked in red., as the university is likely to be seen as a benevolent provider with minimal financial or marketing-driven motives (Nordheim et al., 2019). In this context, university branding refers to the integration of the university’s logo and color scheme into the user interface, creating a visually prominent link between the system and the institution. University branding may also create a sense of familiarity, a factor found to foster interpersonal trust and that might also support trust in AI when users associate it with a familiar institution (Alarcon et al., 2016). Second, the process dimension of the system may be perceived as more trustworthy because it is easily accessible and embedded in the existing organizational framework. Here, features such as the option to contact a human operator for assistance could contribute to users’ trust through perceived openness and accessibility of the system (Lee and See, 2004). Additionally, a customized LLMaaS chatbot may be perceived as more transparent due to context-specific information about the system available on the university’s website. These factors together lead to the hypothesis:
@@
We further hypothesize that higher trust in the organization’s customized LLMaaS chatbot will likely depend on university members’ trust in the university itself. Organizational trust describes the extent to which individuals trust an organization (Pirson and Malhotra, 2011). Prior research suggests that organizational trust can significantly shape users’ trust in the organization’s system (Balayn et al., 2025; Łapińska et al., 2021). In line with this, Nordheim et al. (2019) have demonstrated that customers’ trust in a company’s service chatbot is influenced by their trust in the company itself. A similar mechanism might apply in the context of academic institutions: users’ trust in a university-provided LLMaaS chatbot may be influenced by their trust in the university. We therefore hypothesize:
@@ 2.4.1. Background
When implementing LLMaaS chatbots, universities face the challenge of ensuring that students and staff can use these systems appropriately and securely. A key aspect of this responsibility is addressing the issue of hallucinations. Hallucinations in LLMs describe the generation of content that appears plausible but is factually incorrect or inconsistent with user input (Huang et al., 2025). Hallucinations in human-AI interaction pose a dual challenge: they impair effective collaboration through reduced output quality and undermine users’ trust (Gong et al., 2024; Amaro et al., 2024). In the academic context, hallucinations can pose particular challenges, as they can reduce perceived accuracy and usefulness, increase confusion, and slow task completion (Steinbach et al., 2025). Such issues are compounded by the fact that LLMs have been shown to often fabricate inaccurate scientific references (Aljamaan et al., 2024; Chelli et al., 2024). Despite various efforts and technological advances, hallucinations cannot be completely eliminated from LLMs (Xu et al., 2024; Banerjee et al., 2024). This issue calls for a focus on the human side, with the goal of identifying strategies to help users anticipate, mitigate, and react to hallucinations.
@@ 2.4.2. Customizations and Hypotheses
We consider two customization aspects of the user interface: corporate design and warnings about false information. Research has shown that the presence of logos from reputable companies can create positive associations, enhancing the perceived credibility of mobile systems (Lowry et al., 2005). We argue that branding a chatbot with a university logo (see Figure 2, label 1) leads users to interpret the system as university-approved. This could potentially make users less critical and cautious in their interactions with the university-branded system, particularly with regard to the risk of hallucinations. Another notable customization choice of LLMaaS chatbots is the display of warnings about hallucinations or false information. Research suggests that such warnings can help users detect hallucinations more frequently (Nahar et al., 2024). While ChatGPT’s interface includes a disclaimer below the prompt box stating “ChatGPT can make mistakes. Please check important information”, LLMaaS solutions allow such warnings to be adapted, removed, or altered. In our study, the customized LLMaaS chatbot included a hallucination disclaimer only on the login page, with no warning displayed during the actual chat interaction (see Figure 2, label 3). We argue that the absence of a persistent warning may result in users being less mentally primed to consider hallucinations or to critically evaluate the accuracy of the system’s output. The university-branded interface and the lack of a persistent warning may lead users to a) behave less cautiously with respect to potential hallucinations and b) detect fewer hallucinations overall. Based on this, we propose the following hypotheses:
@@
H4: Users report fewer experienced hallucinations when using the customized LLMaaS chatbot compared to a commercial LLM chatbot.
To assess whether actual differences in hallucination tendencies between the systems might contribute to differences in users’ experienced hallucinations, we evaluate both systems using benchmark datasets (see Section 4.4).
@@ 2.5.1. Background
End users often express concerns about data security and privacy when interacting with LLMs (Ma et al., 2025b; Ali et al., 2025). Reflecting these concerns, recent research has called for privacy-friendly user interfaces of LLM systems (Li et al., 2024; Song et al., 2025; Windl et al., 2025). Such concerns are grounded in tangible risks, including (unauthorized) data retention (Cobbe and Singh, 2021), where providers may retain and use customer data to enhance their models, inference attacks that aim to extract sensitive information from trained models (Kandpal et al., 2024), and the unintended exposure of confidential data during model training or operation (Chen et al., 2024a). To mitigate these risks, technical customization options in LLMaaS offer promising avenues, playing a pivotal role in enhancing users’ perceptions of privacy and security while fostering trust. These options include user-configurable privacy settings that empower users to control how their data is stored and used (Ashley et al., 2002), transparency mechanisms like privacy explanations (Brunotte et al., 2023), which provide clarity about data handling practices, and the implementation of privacy-preserving techniques such as differential privacy (Abadi et al., 2016) during training and fine-tuning processes.
@@ 2.5.2. Customizations and Hypothesis
In the LLMaaS chatbot under study, we identified two system characteristics with potential relevance for users’ privacy perceptions: restricted access to internal networks and the exclusion of chat data from model training (see Figure 1). Beyond these direct technical benefits, we posit that other customization measures – even those without tangible impacts on privacy or security – may foster higher perceived privacy levels. In this context, we again draw on the role of corporate branding in the interface design as a salient customization feature: University branding (see Figure 2, label 1) may enhance users perceptions of data security and privacy, as they may believe that their data will be handled with care and not shared with third parties. Prior research supports this assumption, showing that a strong brand image or high brand credibility can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022). Thus, we propose:
@@ 2.6.1. Background
In this final part, we broaden the perspective to considerations that pertain to the global strategy of an organization, specifically the corporate social responsibility (CSR). CSR refers to situations in which organizations pursue strategies that go beyond business needs and legal regulations but aim at furthering social good (McWilliams and Siegel, 2001). This includes, e. g., improving an organization’s sustainability and environmental performance (McWilliams et al., 2006). We argue that customization of LLMaaS chatbots offers organizations opportunities to align AI deployment with their CSR strategy and corresponding sustainability goals, thereby not only shaping the trust in and perception of the systems but also their environmental impact.
The environmental impact of AI systems and LLMs has become an increasingly critical concern in computing research. Recent work by Rafael and Reis (2024) emphasizes how the invisibility of digital energy consumption often leads to unconscious and potentially wasteful usage behavior. This mirrors findings in energy consumption studies (McCarthy et al., 2023) that identify varying levels of user awareness – from “energy unaware” to “energy aware and active” – highlighting how visibility of resource consumption can transform user behavior. Our research extends these insights and considers the impact of customization choices aimed at promoting sustainability-aware AI usage.
@@ 2.6.2. Customizations and Hypothesis
Making resource consumption visible has shown promising results across different domains. Penkert and Gröschel (2023) demonstrated that transparency in sustainability information significantly influences user choices, while Sanduleac et al. (2017) found that real-time energy data promotes more responsible consumption patterns. LLMaaS chatbots offer customization options to make AI resource limitations explicit through token visibility (see Figure 2, label 2). The examined LLMaaS chatbot’s interface shows the remaining token contingent for the respective language model as a percentage. We argue that customizing token visibility can foster what we term “sustainability-aware AI use” – a more deliberate approach to AI interaction that considers broader environmental impact. Based on these previous findings, we hypothesize:
@@ 3.1. Study Design
We conducted a quantitative, cross-sectional field study. Data was collected through an online survey during spring 2025 (3rd-12th March) at a major public university that had implemented a customized LLMaaS chatbot at the end of 2024. Participants were recruited via the university chatbot landing page, departmental communications, campus flyers, the university news website, and official university social media channels. Participants who reported regular usage of both chatbots answered the same questions for the university’s customized LLMaaS chatbot and for ChatGPT.
@@ 3.2. Description of the University’s Customized LLMaaS Chatbot
The LLMaaS chatbot introduced by the university serves multiple LLMs hosted by different providers. The default model for each chat session, GPT-4o, as well as other OpenAI models, are provided by Microsoft Azure using a datacenter in Sweden. Furthermore, users can choose from a small selection of open-weight models provided by a German AI service center (see Figure 2). The system had 4636 users at the beginning of the data collection period and grew by $20.7\%$ to 5593 users by the end of the period. Customizations regarding the data and model were kept to a minimum, but the university made some adjustments to the user interface and incorporated detailed information and training materials (see Table 2). In contrast to OpenAI’s ChatGPT, the university chatbot supports only text-based interactions and image generation. Other modalities such as audio, document upload, and advanced function calling (e. g., web search) are not supported. Figure 2 shows an anonymized replica of the user interface.
Figure 2. Anonymized replica of the user interface of the university’s customized LLMaaS chatbot. Key elements: (1) university name is displayed; interface follows university’s visual design; (2) remaining monthly token contingent shown in percent; (3) no warning regarding potential hallucinations / incorrect outputs.:
@@
Model fine-tuning and Configuration Temperature1  set to 0 for deterministic outputs. System prompt fixed to: ”You’re a helpful assistant.” for all models, not modifiable by users
Integration and Interoperability Access to multiple models; image generation via DALL-E for OpenAI models
@@
User Interface and Experience Interface customized with corporate design (e. g., university logo); monthly token contingent visualized as a percentage in the user interface; no warning about hallucinations in bottom section; dropdown menu for choosing active model
@@
  • 1
    The temperature parameter in LLMs, ranging from 0 to 1, controls response randomness, with values near 0 favoring deterministic outputs and values near 1 producing diverse, creative outputs
$3.$ Ethical Considerations
The study was designed and conducted with great care and attention to ethical criteria. By involving the a) data protection officer, b) work councils for research and administrative staff, c) center for teaching and learning as well as d) information technology department, the rights and interests of all stakeholders were taken into account in the design and implementation of the study. The approval of all parties was sought prior to implementation. The study was also approved by the local ethics committee. Concerning data collection, storage, and sharing, we followed the guidelines of the university’s IT committee and used committee-approved survey software with university-internal data storage, ensuring that only the project team had access to the raw data. The participants were informed about the content and procedure of the study via standard written instructions in line with the guidelines of the German Research Foundation (Schönbrodt et al., 2017). Participation was voluntary, anonymous, and uncompensated.
@@ 3.4. Survey Structure and Measures
To avoid participant fatigue, all hypothesis-relevant items followed the same response format to minimize cognitive load. We also conducted pilot runs to ensure that average completion time remained under 15 minutes. In addition, only hypothesis-relevant items were mandatory and were placed toward the beginning of the questionnaire to reduce the likelihood of fatigue affecting key measures.
@@
Figure 3. Survey flowchart illustrating participant routing and item blocks based on chatbot usage. Labels at the top of each block indicate the estimated completion time for that section.:
@@
The survey began by capturing usage patterns for both systems: frequency of use (ranging from “never” to “several times a day”) and context of use (6-point scale in response to the item “I tend to use the following AI chatbots for …”, ranging from 1 = “private purposes” to 6 = “study-/work-related purposes”). Next, several aspects of system perception and use were explored: specific use purposes (e. g., information search), perceived usefulness (productivity enhancement, task facilitation), and ease of use (interface clarity, learning curve); the latter two assessed following the Technology Acceptance Model (TAM) (Davis, 1989). For users familiar with either system, we also assessed their awareness of model switching options and use of different large language models available within each platform. For users who have never used or rarely use the university’s LLMaaS, we explored their reasons for non-use or non-regular use.
@@
At the end of the survey, a knowledge test was administered to assess participants’ understanding of the customized LLMaaS chatbot. The test consisted of five items (e. g., “The university chatbot offers models trained by the university” (false)) with response options “true,” “false,” and “I don’t know.” The correct answers of the participants were summed to create a knowledge score ranging from 0 to 5. Participants were debriefed on the following page by being shown the correct answers. At the end of the study, participants could voluntarily report demographic information and were given the opportunity to provide comments in an open text field.
@@
The main part of the survey was structured around the key constructs (see hypotheses). Participants who used both systems responded to the same set of items twice: once for the university’s customized LLMaaS chatbot and once for ChatGPT. Only participants who reported using a system at least once a month were presented with the full set of items about the respective system. Participants were reminded to answer the questionnaire items with regards to their work-related or academic use of the system. We used five-point Likert scales ranging from “strongly disagree” to “strongly agree” for all scales. The questionnaire was available in both German and English, with items from previously published scales translated for the German version. Internal consistencies (Cronbach’s $\alpha$) for the scales are reported in Table B1 and Table B2.
@@
Trust in the two systems was assessed by the facets of global trust, benevolence/purpose, integrity/process, and ability/performance using four items from Wischnewski et al. (2025) (e. g., “I trust [the university chatbot/ChatGPT].”), forming an overall trust measure. We selected this scale for its recent validation, context specificity, and strong fit to human–AI interaction contexts, as well as its theoretical grounding in calibrated trust research (Lee and See, 2004; Parasuraman and Riley, 1997).
@@
Trust in the university was measured with three items following the trust in management scale by Bøe (2018) (e. g., “My university will safeguard my interests.”). We selected this organizational trust scale due to its strong psychometric validation and its applicability to university contexts. In the validation study, it also predicted intention to continue using institutional technology (Bøe, 2018).
@@
We investigated users’ cautious behavior toward LLM hallucinations, including their verification practices and task-specific trust decisions, with three items (e. g., “I check important information from [the university chatbot/ChatGPT] responses through external sources.”). Users also reported their experienced hallucinations with three items (e. g., “I have noticed that [the university chatbot/ChatGPT] makes up facts that do not correspond to reality.”). The items were self-developed, as no existing validated scale captured these aspects of LLM hallucinations. Two of the three items measuring experienced hallucinations build on the first two items of the “AI Hallucination Potential” scale by Christensen et al. (2024), which we adapted to our context. We added a third item based on the definition of hallucinations provided by Huang et al. (2025). The items measuring cautious behavior toward hallucinations capture three key user strategies: deciding on LLM use based on the importance of factual accuracy, verifying outputs externally, and adjusting prompts to improve reliability. These strategies are informed by existing research on LLM hallucination mitigation (Barkley and van der Merwe, 2024; Gordon et al., 2023). The internal consistency of the experienced hallucinations scale was acceptable, while cautious behavior toward hallucinations showed low internal consistency. Item-level analysis indicated that removing individual items did not increase reliability beyond $\alpha=0.41$. We discuss this limitation in Section 5.4. Discriminant validity between the two constructs was supported by low correlations, indicating that cautious behavior and experienced hallucinations represent distinct constructs.
@@
We measured privacy concerns with three items adapted from Hsu and Lin (2018) (e. g., “My decision to use [the university chatbot/ChatGPT] exposes me to privacy risks.”). To increase content validity and contextual relevance for our study, we added two self-developed items. One of them assesses users’ belief in the confidentiality of the system; the other addresses concerns about chat conversations being read by third parties. This addition was motivated by recent research suggesting that such concerns may be particularly relevant for students using chatbots in academic contexts (Chen et al., 2023).

### v3

@@ 2.4.2. Customizations and Hypotheses
We theorize that users report different levels of trust in a customized LLMaaS chatbot compared to a commercially available LLM chatbot. Drawing on the TrAM, we expect customizations of the LLMaaS user interface to serve as a salient trust cue for users, increasing the perceived trustworthiness of the system in two dimensions: (a) purpose and (b) process. (a) The perceived purpose of the LLMaaS system may appear more trustworthy when university branding is incorporated into the front-end of the chatbot (see Figure 2, Label 1) [^1] Figure 2 shows a replica of the LLMaaS chatbot’s user interface, with customizations marked in red., as the university is likely to be seen as a benevolent provider with minimal financial or marketing-driven motives (Nordheim et al., 2019). In this context, university branding refers to the integration of the university’s logo and color scheme into the user interface, creating a visually prominent link between the system and the institution. The incorporation of university branding in the front-end may also create a sense of familiarity, a factor found to foster interpersonal trust and that might also support trust in AI when users associate it with a familiar institution (Alarcon et al., 2016). (b) The process dimension of the system may be perceived as more trustworthy because it is easily accessible and embedded in the existing organizational framework. Here, features such as the option to contact a human operator for assistance could contribute to users’ trust through perceived openness and accessibility of the system (Lee and See, 2004). Additionally, the customized LLMaaS chatbot may be perceived as more transparent due to context-specific information about the system available on the university’s website. These factors together lead to the hypothesis:
@@
We further hypothesize that higher trust in the organization’s customized LLMaaS chatbot will likely depend on university members’ trust in the university itself. In reference to the TrAM, we consider organizational trust to be an external, moderating factor that has an impact on the association between perceived system characteristics and reported trust in a system. Organizational trust describes the extent to which individuals trust an organization (Pirson and Malhotra, 2011). Prior research suggests that organizational trust can significantly shape users’ trust in the organization’s system (Balayn et al., 2025; Łapińska et al., 2021). In line with this, Nordheim et al. (2019) have demonstrated that customers’ trust in a company’s service chatbot is influenced by their trust in the company itself. A similar mechanism might apply in the context of academic institutions: users’ trust in a university-provided LLMaaS chatbot may be influenced by their trust in the university. We therefore hypothesize:
@@ 2.5.1. Background
Hallucinations in LLMs refer to the generation of content that appears plausible but is factually incorrect or inconsistent with user input (Huang et al., 2025). Hallucinations in human-AI interaction pose a dual challenge: they impair effective collaboration by reducing output quality and undermining users’ trust (Gong et al., 2024; Amaro et al., 2024). In the academic context, hallucinations can pose particular challenges, as they can reduce perceived accuracy and usefulness, increase confusion, and prolong task completion (Steinbach et al., 2025). Such issues are compounded by the fact that LLMs have been shown to often fabricate inaccurate scientific references (Aljamaan et al., 2024; Chelli et al., 2024).
@@ 2.5.2. Customizations and Hypotheses
We expect there to be differences with regard to users’ perception of and cautious behavior towards hallucinations and ground this expectation in two customization aspects: branding through corporate design and warnings about false information. Thus, with reference to the TrAM, we conceptualize branding with corporate design and warnings about false information as system cues and consider users’ perception of and cautious behavior towards hallucinations as important ’outcome’ perception dimensions. Research has shown that the presence of logos from reputable companies can create positive associations, enhancing the perceived credibility of mobile systems (Lowry et al., 2005). We argue that branding a chatbot with a university logo (see Figure 2, label 1) leads users to interpret the system as university-approved. This could potentially make users less critical and cautious in their interactions with the university-branded system, particularly with regard to the risk of hallucinations. Another notable customization choice of LLMaaS chatbots is the display of warnings about hallucinations or false information. Research suggests that such warnings can help users detect hallucinations more frequently (Nahar et al., 2024). While ChatGPT’s interface includes a disclaimer below the prompt box stating “ChatGPT can make mistakes. Please check important information”, LLMaaS solutions allow such warnings to be adapted, removed, or altered. In our study, the customized LLMaaS chatbot included a hallucination disclaimer only on the login page, with no warning displayed during the actual chat interaction (see Figure 2, label 3). We argue that the absence of a persistent warning may result in users being less mentally primed to consider hallucinations or to critically evaluate the accuracy of the system’s output. The university-branded interface and the lack of a persistent warning may lead users to a) behave less cautiously with respect to potential hallucinations and b) detect fewer hallucinations overall. Based on this, we propose the following hypotheses:
@@
  • H4: Users report fewer perceived hallucinations when using the customized LLMaaS chatbot compared to a commercial LLM chatbot.
To assess whether actual differences in hallucination tendencies between the systems might contribute to differences in users’ perceived hallucinations, we evaluate both systems using benchmark datasets (see Section 4.5).
Figure 2. Anonymized replica of the user interface of the university’s customized LLMaaS chatbot. Key elements: (1) university name is displayed; interface follows university’s visual design; (2) remaining monthly token contingent shown in percent; (3) no warning regarding potential hallucinations / incorrect outputs.
Screenshot of the anonymized replica of the university’s customized LLMaaS chatbot interface. The top navigation bar is blue with four buttons on the right: Chat, Models, Settings, and Logout. On the left, the label “UniGPT” is displayed, styled to match the university’s visual identity (1). Beneath the navigation bar is the chat area. On the left panel, a chat history lists a single entry “Hello UniGPT!”. The main panel shows a conversation with a greeting from the chatbot (“Hello! How can I assist you today?”). A drop-down menu labeled “Models” is open, listing model options such as GPT-4o, GPT-3.5 Turbo, GPT-4, Meta LLaMa, Mistral Large Instruct, NVIDIA LLaMa, Codestarl, and Teuken 7B. On the right side of the main panel, a status element (2) displays the remaining monthly token contingent as a percentage (“Tokens available: 70.15%”). Below the chat window is a text entry field with a send button on the right. Area (3) is an empty space, highlighting the absence of any visual warning about potential hallucinations or incorrect outputs in the interface.
@@ 2.6.1. Background
End users often express concerns about data security and privacy when interacting with LLMs (Ma et al., 2025; Ali et al., 2025). Reflecting these concerns, recent research has called for privacy-friendly user interfaces of LLM systems (Li et al., 2024; Song et al., 2025; Windl et al., 2025). Such concerns are grounded in tangible risks, including (unauthorized) data retention (Cobbe and Singh, 2021), where providers may retain and use customer data to enhance their models, inference attacks that aim to extract sensitive information from trained models (Kandpal et al., 2024), and the unintended exposure of confidential data during model training or operation (Chen et al., 2024). To mitigate these risks, technical customization options in LLMaaS offer promising avenues, playing a pivotal role in enhancing users’ perceptions of privacy and security while fostering trust. These options include user-configurable privacy settings that empower users to control how their data is stored and used (Ashley et al., 2002), transparency mechanisms like privacy explanations (Brunotte et al., 2023), which provide clarity about data handling practices, and the implementation of privacy-preserving techniques such as differential privacy (Abadi et al., 2016) during training and fine-tuning processes.
@@ 2.6.2. Customizations and Hypothesis
In the LLMaaS chatbot under study, we identified two system characteristics with potential relevance for users’ privacy perceptions: restricted access to internal networks and the exclusion of chat data from model training (see Figure 1). In reference to the TrAM, these represent actual system characteristics, which users might be able to perceive/ inform themselves about. Beyond these direct technical benefits, we posit that other customization measures – even those without tangible impacts on privacy or security – may foster higher perceived privacy levels. In regard to the TrAM, we conceptualize those as perceived system characteristics that relate to users’ privacy concerns but not to the actual systems’ privacy settings. In this context, we again draw on the role of corporate branding in the interface design as a salient customization feature: University branding (see Figure 2, label 1) may enhance users perceptions of data security and privacy, as they may believe that their data will be handled with care and not shared with third parties. Prior research supports this assumption, showing that a strong brand image or high brand credibility can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022). Thus, we propose:
@@ 2.7.1. Background
In this final part, we broaden the perspective to considerations that pertain to the global strategy of an organization, specifically the corporate social responsibility (CSR). CSR refers to situations in which organizations pursue strategies that go beyond business needs and legal regulations but aim at furthering social good (McWilliams and Siegel, 2001), e. g., improving an organization’s environmental performance (McWilliams et al., 2006). We argue that customization of LLMaaS chatbots offers organizations opportunities to align AI deployment with their CSR strategy and corresponding sustainability goals, thereby not only shaping the trust in and perception of the systems but also their environmental impact.
The environmental impact of AI systems and LLMs has become an increasingly critical concern in computing research. Recent work by Rafael and Reis (2024) emphasizes how the invisibility of digital energy consumption often leads to unconscious and potentially wasteful usage behavior. Our research builds on these insights and considers the role of customization choices aimed at promoting sustainability-aware AI usage.
@@ 2.7.2. Customizations and Hypothesis
Making resource consumption visible has shown promising results across different domains. Penkert and Gröschel (2023) demonstrated that transparency in sustainability information significantly influences user choices, while Sanduleac et al. (2017) found that real-time energy data promotes more responsible consumption patterns. LLMaaS chatbots offer customization options to make AI resource limitations explicit through token visibility (see Figure 2, label 2). We argue that customizing token visibility can foster what we term “sustainability-aware AI use” – a more deliberate approach to AI interaction that considers broader environmental impact. In reference to the TrAM, the displayed tokens represent perceived system characteristics that are salient to the user. We theorize that those might lead to differences in self-reported sustainability aware AI use. Based on these previous findings, we hypothesize:
@@ 3.1. Study Design
We conducted a quantitative, cross-sectional field study. Data was collected through an online survey (3rd March - 12th May 2025) at a major public university that had implemented a customized LLMaaS chatbot at the end of 2024. Participants who reported regular usage of both chatbots answered the same questions for the university’s customized LLMaaS chatbot and for ChatGPT.
@@ 3.2. Description of the University’s Customized LLMaaS Chatbot
The LLMaaS chatbot introduced by the university serves multiple LLMs hosted by different providers. The default model for each chat session, GPT-4o, as well as other OpenAI models, are provided by Microsoft Azure using a datacenter in Sweden. Furthermore, users can choose from a small selection of open-weight models provided by a German AI service center (see Figure 2). Thus, only application layers (i) front-end and (ii) back-end are under the university’s control. The system had 4636 users at the beginning of the data collection period and grew by $20.7\%$ to 5593 users by the end of the period. Customizations regarding the back-end were kept to a minimum, but the university made some adjustments at the front-end, specifically to the user interface and incorporated detailed information (see Table 1). In contrast to OpenAI’s ChatGPT, the university chatbot supports only text-based interactions and image generation. Figure 2 shows an anonymized replica of the user interface.
@@
Model Fine-tuning and Configuration No fine-tuning or other model customizations. Temperature1 set to 0 for deterministic outputs. System prompt fixed to: ”You’re a helpful assistant.” for all models, not modifiable by users
Integration and Interoperability Access to multiple models from different LLMaaS providers; image generation via DALL-E for OpenAI models
@@
User Interface and Experience Custom-built interface following corporate design (e. g., university logo); monthly token contingent visualized as a percentage in the user interface; no warning about hallucinations at the bottom section; dropdown menu for choosing active model
@@
1 The temperature parameter in LLMs, ranging from 0 to 1, controls response randomness, with values near 0 favoring deterministic outputs and values near 1 producing diverse, creative outputs.
3.3. Ethical Considerations
By involving the a) data protection officer, b) work councils, c) center for teaching and learning and d) information technology department, the rights and interests of all stakeholders were taken into account in the design of the study. The approval of all parties and the local ethics committee was sought prior to implementation. Concerning data collection, storage, and sharing, we followed the guidelines of the university’s IT committee and used committee-approved survey software with university-internal data storage, ensuring that only the project team had access to the raw data. The participants were informed about the content and procedure of the study via standard written instructions (Schönbrodt et al., 2017). Participation was voluntary, anonymous, and uncompensated.
@@ 3.4. Survey Structure and Measures
All hypothesis-relevant items followed the same response format to minimize cognitive load and participant fatigue. We also conducted pilot runs to ensure that the average completion time remained under 15 minutes. Only hypothesis-relevant items were mandatory and were placed toward the beginning of the questionnaire to reduce the likelihood of fatigue affecting key measures.
@@
Figure 3. Survey flowchart illustrating participant routing and item blocks based on chatbot usage. Labels at the top of each block indicate the estimated completion time for that section.
Flowchart with six labeled blocks, connected by directional arrows, representing the survey structure for different participant groups. The sequence begins with an introductory module (“Introduction, survey explanation, informed consent”, 1 min). Next, all participants complete a block with two items on chatbot usage frequency and types of AI chatbot use (1 min). The flow then branches based on chatbot usage: participants complete either (a) a 32-item module for users of the university’s customized chatbot (5 min), or (b) a four-item module for non-users of this chatbot (2 min). Participants who report using ChatGPT additionally complete a 30-item module (5 min). All participants finish with a final 14-item module that covers trust in the university, general hallucination awareness, the knowledge test, and demographics (3 min).
@@
The survey began by capturing usage patterns for both systems: frequency of use (ranging from “never” to “several times a day”) and context of use (6-point scale in response to the item “I tend to use the following AI chatbots for …”, ranging from 1 = “private purposes” to 6 = “study-/work-related purposes”). Next, several aspects of system perception and use were explored: specific use purposes (e. g., information search), perceived usefulness (productivity enhancement, task facilitation), and ease of use (interface clarity, learning curve); the latter two were assessed following the Technology Acceptance Model (TAM) (Davis, 1989). For users familiar with either system, we also assessed their awareness of model switching options and use of different models. For users who have never or rarely used the university’s LLMaaS, we explored their reasons for non-use.
@@
At the end of the survey, a knowledge test was administered to assess participants’ understanding of the customized LLMaaS chatbot. The test consisted of five items (e. g., “The university chatbot offers models trained by the university” (false)) with response options “true”, “false”, and “I don’t know”. Correct answers were summed to create a knowledge score ranging from 0 to 5. Participants were debriefed on the following page about the correct answers. At the end of the study, participants could voluntarily report demographic information.
@@
The main part of the survey was structured around the key constructs. Participants who used both systems responded to the same set of items twice: once for the university’s customized LLMaaS chatbot and once for ChatGPT. Only participants who reported using a system at least once a month were presented with the full set of items. Participants were reminded to answer the questionnaire items with regards to their work-related or academic use of the system. We used five-point Likert scales ranging from “strongly disagree” to “strongly agree” for all scales. The questionnaire was available in German and English, with items from previously published scales translated for the German version. Internal consistencies (Cronbach’s $\alpha$) for the scales are reported in Table B1 and Table B2.
@@
Trust in the systems was assessed by the facets of global trust, benevolence/purpose, integrity/process, and ability/performance using four items from Wischnewski et al. (2025) (e. g., “I trust [the university chatbot/ChatGPT].”), forming an overall trust measure. We selected this scale for its recent validation, context specificity, and strong fit to human–AI interaction contexts, as well as its theoretical grounding in calibrated trust research (Lee and See, 2004; Parasuraman and Riley, 1997).
@@
Trust in the university was measured with three items following the trust in management scale by Bøe (2018) (e. g., “My university will safeguard my interests.”). We selected this scale due to its strong psychometric validation and its applicability to university contexts. In the validation study, it also predicted intention to continue using institutional technology (Bøe, 2018).
@@
We investigated users’ cautious behavior toward LLM hallucinations, including their verification practices and task-specific trust decisions, with three items (e. g., “I check important information from [the university chatbot/ChatGPT] responses through external sources.”). Users also reported their perceived hallucinations with three items (e. g., “I have noticed that [the university chatbot/ChatGPT] makes up facts that do not correspond to reality.”). The items were self-developed, as no existing validated scale captured these aspects of LLM hallucinations. Two of the three items measuring perceived hallucinations build on the first two items of the “AI Hallucination Potential” scale by Christensen et al. (2024), which we adapted to our context. We added a third item based on the definition of hallucinations provided by Huang et al. (2025). The items measuring cautious behavior toward hallucinations capture three key user strategies: deciding on LLM use based on the importance of factual accuracy, verifying outputs externally, and adjusting prompts to improve reliability. These strategies are informed by existing research on LLM hallucination mitigation (Barkley and van der Merwe, 2024; Gordon et al., 2023). The internal consistency of the perceived hallucinations scale was acceptable, while cautious behavior toward hallucinations showed low internal consistency. Item-level analysis indicated that removing individual items did not increase reliability beyond $\alpha=0.41$. We discuss this limitation in Section 5.3. Discriminant validity between the two constructs was supported by low correlations, indicating that they represent distinct constructs.
@@
We measured privacy concerns with three items adapted from Hsu and Lin (2018) (e. g., “My decision to use [the university chatbot/ChatGPT] exposes me to privacy risks.”). To increase content validity and contextual relevance for our study, we added two self-developed items. One assessing users’ belief in the confidentiality of the system; the other concerns about chat conversations being read by third parties. This addition was motivated by recent research suggesting that such concerns may be particularly relevant for students in academic contexts (Chen et al., 2023).